# Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation

### v1

@@ 4.2 Results and Discussion
Table 4: Performance of different hop numbers. The best result is shown in bold.
@@
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
GFM-RAG 78.22 69.40 32.78 53.78 34.63 7.02 40.35 30.25 1.45
RAPTOR 77.39 71.25 44.86 53.20 38.35 8.57 42.82 33.14 2.76
GraphRAG - - - 54.99 39.91 9.79 45.18 34.38 3.22
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.2 35.76 3.78
QSGNN(Ours) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
Table 5: Retrieval efficiency on MuSiQue. Time means the average retrieval time of each query.
Method NV-Embed-v2 GFM-RAG RAPTOR HippoRAG2 QSGNN
Recall@5 68.19 67.19 69.55 72.29 75.23
Time(s) 0.036 0.086 0.376 0.317 0.118
Efficiency. We evaluate the efficiency of retrieval across different methods, which is summarized in Table 5. Graph search based methods (HippoRAG2 and RAPTOR) exhibit higher time cost because the graph search strategies are expensive. GNN based methods (QSGNN and GFM-RAG) cost less time since message passing is an efficient way to aggregate graph information  (Hamilton et al., 2017). QSGNN shows more retrieval time than GFM-RAG due to the complexity of Multi-L KG.
Ablation Study. To evaluate the influence of different components, we conduct ablation studies and report the results in Table 6. Our key findings are: i) The model’s performance drops significantly without entities, as they form the basic semantic relationships in Multi-L KGs. Removing entities causes substantial knowledge loss. ii) No intra-level message passing severely impacts performance because QSGNN loses the ability to directly capture Multi-L KG structures but have to be bridged by chunk nodes. iii) The absence of query alignment has a particularly negative effect on high-hop questions, where the influence of noise will be amplified as the hop number increases. iv) Removing chunk nodes shows relatively less impact, as document nodes can provide global contextual information to compensate for their absence. More results are shown in Appendix  A.6.
Table 6: Ablation study of QSGNN. “w/o q_attn” means no alignment with query. “w/o intra” means no intra-level message passing. “w/o entity” and “w/o chunk” mean no entity or chunk in Multi-L KG. No inter-level message passing and No document are not reported here since we need retrieve documents.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
Other Discussion. For the page limitation, we discuss the hyperparameter influence of QSGNN in Appendix  A.7. We conduct the good case study in Appendix  A.8. The bad case study and the discussion of limitations for QSGNN can be found in Appendix  A.9.
5 Conclusion
@@
In this paper, we propose a novel graph representation learning framework for multi-hop questions retrieval. We first design a Multi-information Level Knowledge Graph (Multi-L KG) to model the multi-granular relationships among documents. Then we introduce Query-Specific Graph Neural Network (QSGNN). The QSGNN aggregates information from multi-levels, together with query alignment, to ensure comprehensive query-specific representations while filtering out the noise. The pre-training strategy further enhance the ability of QSGNN. Extensive experiments shows the effectiveness of our method in multi-hop questions.
Ethics Statement. We affirm that our research adheres to the highest ethical standards in the conduct of scientific research. This work does not involve human subjects, animal testing, harmful insights, discrimination, sensitive personal data, or any legal compliance problem. All experiments were conducted using publicly available datasets or Large Language Model (LLM) tools, and we have taken steps to ensure the reproducibility and transparency of our results. We have considered potential societal impacts of our research and believe that the proposed methods and findings contribute positively to the field of LLMs and its applications.
Reproducibility Statement. To ensure the reproducibility of our work, we provide the following statement. All code implementations of our proposed method are provided by an anonymous link, and we will make them publicly open source upon the acceptance of this paper. Additionally, all the large language models (LLMs) used in this paper are open source, as detailed in prior work  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025). Regarding the data used in our experiments, we follow the standard protocols established by open source works  (Gutiérrez et al., 2025). All the data processing details are presented in Appendix A.2. For better reproducibility, we also include example data in our code repository, which will be released publicly after the acceptance of the paper.
@@ Appendix A Appendix
This supplementary material provides additional details on the proposed method and experimental results that could not be included in the main manuscript due to page limitations. Specifically, this appendix is organized as follows.
  • Sec. A.1 provides the use of Large Language Models (LLMs).
  • Sec. A.2 provides more details on the dataset settings and data processing.
  • Sec. A.3 outlines the implementation of baseline models and their experimental settings.
  • Sec. A.4 presents the QA performance on GPT-4o-mini.
  • Sec. A.5 provides the multi-hop performance on 2WikimultihopQA(2Wiki) dataset.
  • Sec. A.6 provides detailed ablation study for our method.
  • Sec. A.7 presents hyperparameter study for our method.
  • Sec. A.8 discusses three good cases of our method compared to HippoRAG2
  • Sec. A.9 presents three bad cases of our method and discusses the limitation and possible solution as future work.
  • Sec. A.10 presents the prompts used for OpenIE.
  • Sec. A.11 presents the prompts used for generating pre-training data.
A.1 The use of Large Language Models
We present the statement regarding the usage of Large Language Models (LLMs) in our paper. We assert that there is no involvement of LLMs in the innovation, implementation, or paper writing of our method. However, we still use LLMs in two ways:
i) Our proposed method is about the Retrieval-augmented generation (RAG) for LLMs. Therefore, in this paper, we use LLMs as the QA performer and the OpenIE tool to evaluate the effectiveness of our method, similar to other LLM-related works  (Guo et al., 2024; Jimenez Gutierrez et al., 2024; Luo et al., 2025).
ii) We use LLMs to help us correct some grammatical errors in our paper writing.
A.2 Dataset Settings
Table 7: Dataset statistics
@@ A.2 Dataset Settings
#Eval QA 1,000 1000 1000
Our dataset is constructed based on the benchmark datasets used by prior works (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025; Luo et al., 2025). Here we present the construction details.
Dataset Construction. The evaluation set is the same as HippoRAG2(Gutiérrez et al., 2025), which is the state-of-the-art baseline. And the corpora is expanded. Since our method need extra QA pairs for fine-tuning, we randomly sample extra samples from each original dataset and we make sure there is no overlap with the evaluation set. For the MuSiQue, the sample should have at least 60% supported documents contained in existing corpora. For the 2Wiki, the sample should have at least 60% overlap with the corpora. For the HotpotQA, the overlap ratio is 50%. Within these extra samples, we randomly select 1000 questions for fine-tuning and 225 questions for testing. The corpora of each dataset will be expanded to cover the new samples related documents(supported documents and distracted documents).
Pre-training Data Construction. We generate synthesized pre-training data for each dataset. For the MuSiQue the one-hop QA pairs are 91,621 and the two-hop QA pairs are 58,923. For the 2Wiki, the one-hop QA pairs are 52,122 and the two-hop QA pairs are 12,097. For the HotpotQA, the one-hop QA pairs are 73,128 and the two-hop QA pairs are 20,619. We combine and shuffle all these QA pairs of each dataset as pre-training data.
The statistics of these datasets are shown in Table  7.
A.3 Baseline Settings
Implementation. The BM25 retrieval algorithm is implemented as BM25S (Lù, 2024). The Contriever  (Izacard et al., 2021) is implemented with the official code acting as the retrieval server in IRCoT  (Trivedi et al., 2022b). As for the text embedding based model, we use the NV-Embed-v2-7B and GTE-Qwen2-7B-Instruct from Huggingface  (Wolf et al., 2019). For the all the GNN based methods and graph search based methods we use their official implementations.
Experimental Settings. For all the baselines, we conduct experiment under the official settings or prior works experimental suggestions. Specifically, their settings are described as follows:
i) For BM25, we retrieve 5 most relevant documents based on the query terms frequency appearing in each document.
ii) For Contriever+IRCoT, we follow the suggestion of HippoRAG2  (Gutiérrez et al., 2025) that we retrieve the top 10 documents for each step, the maximum step number for MuSiQue is set as 4 and the maximum step number for other datasets is set as 2.
iii) For text embedding based methods, embeddings are calculated for all the documents within our datasets, we retrieve the 5 most relevant documents as  (Jimenez Gutierrez et al., 2024).
iv) For GFM-RAG, we use the official model implementation where we pre-train and fine-tune the GNN on our corpora and all the hyperparameters are set as  (Luo et al., 2025) suggests. We retrieve the 5 most relevant documents.
v) For GNN-RAG, we use ReaRev  (Mavromatis & Karypis, 2022) as GNN reasoning, it samples dense subgraph by PageRank  (Andersen et al., 2006) on our datasets, all the settings are the same as the official settings  (Mavromatis & Karypis, 2024).
vi) For GraphRAG and LightRAG, the implementations are based on the official codes, and the hyperparameters are set as HippoRAG2  (Gutiérrez et al., 2025) suggests.
vii) For RAPTOR, HippoRAG and HippoRAG2, the settings are set the same as their official implementations.
For all the methods, we use LLM (Llama-3.3-70B-Instruct  (AI@Meta., 2024) and GPT-4o-mini  (OpenAI., 2024)) for QA task.
A.4 QA Performance on GPT-4o-mini
@@
Table 8: QA performance on GPT-4o-mini. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line.
MuSiQue 2Wiki HotpotQA Average
Method EM F1 EM F1 EM F1 EM F1
No Retrieval 10.81 21.39 31.83 36.51 27.44 41.47 23.36 33.12
BM25 12.54 23.88 32.25 40.30 34.50 52.62 26.43 38.93
Contriever+IRCoT 13.90 25.28 29.44 34.66 30.17 45.55 24.50 35.16
GTE-Qwen2-7B 31.15 39.88 50.75 52.37 52.19 66.17 44.70 52.81
NV-Embed-v2 28.07 37.15 49.22 51.51 54.85 68.43 44.05 52.36
GNN-RAG 23.02 30.12 34.52 48.25 49.12 58.99 35.55 45.79
GFM-RAG 28.30 36.81 43.79 53.26 51.86 64.73 41.32 51.6
RAPTOR 32.93 41.04 45.43 56.24 55.19 68.16 44.52 55.15
GraphRAG 33.31 40.13 50.99 58.08 56.45 69.24 46.92 55.82
HippoRAG2 33.12 40.30 51.49 60.18 56.39 70.53 47.00 57.00
QSGNN(Ours) 34.27 41.78 53.52 61.63 56.04 69.73 47.94 57.71
@@
QA Task on GPT-4o-mini. Table 8 presents QA performance using GPT-4o-mini. We do not report the performance of LightRAG as it is unsuitable for multi-hop QA tasks. The results show that QSGNN achieves competitive EM and F1 scores, demonstrating its compatibility with different LLM backbones. No Retrieval gets the lowest EM and F1 score due to the limited knowledge of GPT-4o-mini. Notably, all methods show reduced performance compared to Llama-3-70B-Instruct(see Table  3), primarily due to the intrinsic knowledge gap between GPT-4o-mini and Llama-3.3-70B-Instruct. Other observations align with the report of Table  3.
A.5 Multi-hop Performance on 2Wiki
Table 9: Performance of different hop numbers on 2Wiki. The best result is shown in bold.
2Wiki(Recall@5) 2Wiki(F1) 2Wiki(EM)
Method 2-hop 4-hop 2-hop 4-hop 2-hop 4-hop
NV-Embed-v2 81.40 63.52 63.21 42.19 59.28 23.35
GFM-RAG 83.37 56.09 64.36 33.11 54.44 18.82
RAPTOR 85.91 64.75 66.13 40.14 53.71 20.12
GraphRAG - - 67.99 39.52 57.24 25.97
HippoRAG2 93.15 65.44 70.50 42.26 61.53 29.32
QSGNN 92.10 74.88 69.72 50.82 60.04 40.28
The multi-hop performance on 2Wiki dataset is presented in Table  9. We find that QSGNN can achieve +14.4%, +20.2% and +37.3% on Recall@5, F1 and EM respectively. For the other baselines, we have the same observations as Table  4.
A.6 Ablation Study
Table 10: Ablation study of QSGNN. “chunk” means performing QA task with chunk. “w/o inter” means no inter-level message passing. “w/o doc” means no document in Multi-L KG. “w/o pt” means no pre-training. “w/o ft” means no fine-tuning.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
QSGNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
QSGNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
QSGNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
QSGNN(chunk) - - - 40.18 37.55 13.26 30.25 29.41 8.71
QSGNN(w/o inter) - - - 40.27 35.98 4.24 28.36 19.19 2.72
QSGNN(w/o doc) - - - 36.54 32.49 8.01 26.19 26.57 5.02
QSGNN(w/o pt) 63.66 53.93 36.83 35.64 28.14 5.15 24.62 19.28 2.53
QSGNN(w/o ft) 77.19 74.69 62.57 51.08 46.42 17.12 43.58 34.23 10.48
We present the ablation study of QSGNN on Table  10. As for the chunk retrieval (QSGNN + chunk, w/o inter or w/o doc), we retrieve the top 10 most relevant chunks for QA and we only report the EM and F1 score since we have no ground truth for chunk retrieval. We find that QSGNN(chunk) does not perform as well as QSGNN. It can be attributed to two factors: i) QSGNN is not directly trained on chunk labels. ii) Document generally contains more information than chunk. No inter-level message passing will not update the representations thus have no difference with the embeddings baselines, whose performance decrease a lot when hop number increases. No document lacks global context information, so it performs worse than QSGNN(chunk). In addition, QSGNN(w/o ft) can achieve competitive performance in 2-hop questions but fail to perform as well as QSGNN, since we pre-train QSGNN only on 1,2-hop questions. QSGNN(w/o pt) struggles to perform well because it can not be fitted well on only small number of training pairs.
A.7 Hyperparameter Discussion
Here we discussion some of important hyperparameters, including the receptive field (QSGNN layer number), negative sampling number, pre-training data scale and information bottleneck dimension.
Receptive Field Discussion. We discuss the relationships between GNN layers and the multi-hop performance, the results are shown in Table  11. We find that 1-layer QSGNN (one intra-level + one inter-level) have limited performance and the gap between 2-layer QSGNN becomes bigger as the hop number increase, it is because 1-layer QSGNN can only aggregate 2-hop information. The receptive field limits the potential of QSGNN. The 2-layer QSGNN achieves the best performance among 2,3,4 hop questions because it can aggregate more useful information. However, as the layer number increases(4 layers), the performance will suffer a lot because of the over-smoothing problem of GNN  (Rusch et al., 2023; Chen et al., 2020a).
Table 11: The influence of GNN layer number.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN(1 layer) 73.81 69.14 55.13 46.25 41.86 11.46 38.29 29.60 7.83
QSGNN(2 layers) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(3 layers) 76.35 74.55 61.66 51.34 45.24 16.45 42.11 34.50 9.13
QSGNN(4 layers) 62.60 58.29 35.37 38.93 29.15 4.61 26.68 18.68 2.03
Negative Sampling Number. Table 12 shows the impact of negative sampling number on model’s performance. According to the results, we find that low sampling number (Neg_N = 1) has limited performance due to inadequate discrimination between target documents and distractions. High sampling number (Neg_N=50, 100) shows limited benefits, because excessive easy negatives fail to provide meaningful guidance for QSGNN training.
Table 12: The influence of negative sampling number.
MuSiQue(Recall@5) MuSiQue(F1) MuSiQue(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.20 35.76 3.78
QSGNN(Neg_N=1) 68.67 60.27 50.82 43.72 38.25 5.18 34.87 27.44 2.10
QSGNN(Neg_N=10) 77.97 76.16 62.14 51.03 46.68 17.62 42.28 38.54 10.76
QSGNN(Neg_N=30) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
QSGNN(Neg_N=50) 77.18 76.69 63.51 51.34 46.99 18.36 44.07 38.02 11.65
QSGNN(Neg_N=100) 77.09 76.16 63.85 52.95 46.37 18.22 43.16 37.52 11.53
Figure 3: Influence of pre-training scale on various model dimension.
Pre-training Scale and Information Dimension. We pre-train QSGNN on the sythesized QA pairs from 0 to 150k across information dimension from 32 to 512 respectively, the results are shown in Figure  3. We conduct experiments on the MuSiQue dataset. As for the pre-training scale, the results show that insufficient pre-training data leads to sub-optimal performance across all the dimensions. As the amount of pre-training data increases, the performance of QSGNN gets better, although the marginal improvement decreases. As for the information dimension, the results indicate that increasing the information dimension improves the performance of QSGNN, as larger dimensions enhance model expressiveness through increased model scale. This observation is consistent with the model scaling analysis presented in GFM-RAG  (Luo et al., 2025). We believe that further improvements can be achieved by combining a larger model with a substantial amount of pre-training data. In this paper, we set the dimension as 128 considering the balance between performance and computational limitation.
A.8 Good Case Studies

### v2

@@ 4.1.4. Implementation Details
Table 3. Performance under different hop numbers. The best result is shown in bold.
MuSiQue(Rec@5) MuSiQue(F1) MuSiQue(EM) 2Wiki(Rec@5) 2Wiki(F1) 2Wiki(EM)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop 2-hop 4-hop 2-hop 4-hop 2-hop 4-hop
NV-Embed-v2 76.86 70.71 39.67 51.09 37.98 9.24 41.67 32.27 3.07 81.40 63.52 63.21 42.19 59.28 23.35
KiRAG 75.42 64.57 26.31 49.24 33.21 6.88 38.93 28.28 1.22 84.58 48.19 64.89 29.88 55.08 14.76
G-retriever 79.16 67.41 29.86 54.29 31.06 6.60 41.02 29.38 1.66 82.36 54.19 63.05 30.67 52.48 15.63
GFM-RAG 78.22 69.40 32.78 53.78 34.63 7.02 40.35 30.25 1.45 83.37 56.09 64.36 33.11 54.44 18.82
GraphRAG - - - 54.99 39.91 9.79 45.18 34.38 3.22 - - 67.99 39.52 57.24 25.97
RAPTOR 77.39 71.25 44.86 53.20 38.35 8.57 42.82 33.14 2.76 85.91 64.75 66.13 40.14 53.71 20.12
ToG2 76.82 69.71 40.81 53.85 37.54 8.23 43.18 33.66 2.43 84.50 63.09 66.23 39.51 52.63 20.69
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24 46.2 35.76 3.78 93.15 65.44 70.50 42.26 61.53 29.32
Quest-GNN(Ours) 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50 92.10 74.88 69.72 50.82 60.04 40.28
4.2. Retrieval and QA Performance
Table 2 shows the retrieval and QA performance of all the methods. We have following observations:
  • No Retrieval achieves low EM and F1 scores, demonstrating the importance of external knowledge for the LLM’s response.
  • Quest-GNN outperforms all baseline methods, especially compared to other GNN based methods. Recall@5 achieves +11.9%, +12.9% and +5.4% improvements and F1 score gets +15.0%, +11.7% and +9.7% improvements on MuSiQue, 2Wiki and HotpotQA respectively, showing its effectiveness on multi-hop questions.
  • The classic retrieval algorithms (BM25/Contriever) and text embedding baselines have limited performance as they fail to capture document relationships which is crucial for answering multi-hop questions.
  • Iterative retrieval based methods show no benefit to multi-hop questions. On the one hand, they only model the question structure but ignore the multi-level relationships between documents. On the other hand, they significantly suffer from intermediate inaccuracies as the suboptimal results may deflect the whole reasoning process.
  • Although GNN based methods are designed to capture multi-hop relationships, their message passing mechanisms may inadvertently aggregate irrelevant information, which degrades retrieval effectiveness and LLM’s response.
  • Graph search based methods generally perform better than other baselines, but their performance is still constrained by inherent limitations. The heuristic exploration rules often overlook multi-level information and may favor specific path patterns, resulting in data bias. For instance, the performance of RAPTOR varies a lot since it leverages Gaussian Mixture Model to detect clusters in KGs, which is highly biased on data. Notably, LightRAG struggles to perform well on these datasets, which is consistent with other reports  (Gutiérrez et al., 2025; Luo et al., 2025). The main reason is that LightRAG is incompatible with multi-hop retrieval workflows as discussed in  (Luo et al., 2025).
We report the QA performance using GPT-4o-mini in Appendix  A.3.
4.3. Multi-hop Performance
In order to analyze the detailed performance on multi-hop questions, we evaluate the performance of the competitive baselines on different hop numbers. The results are shown in Table 3 and we visualize the performance on MuSiQue as Figure  1. We make the following observations:
  • All the methods exhibit performance decline as hop number increases, because of the exponentially growing search space and noise accumulation.
  • Iterative retrieval based methods underperform in high-hop scenarios due to: i) the negelect of complex relationships between documents, ii) accumulated errors from intermediate results.
  • Compared to graph search based methods, existing GNN based methods have more performance decrease in high-hop questions because the message passing aggregates the information of irrelevant nodes without question alignment.
  • As for Quest-GNN, it achieves competitive performance in 2-hop questions compared to other baselines and the improvements become noticeable as the hop number increases. For example, Quest-GNN gets +33.8%, +85.8% and +231% improvements respectively in terms of Recall@5, F1 and EM on 4-hop questions of MuSiQue. This advantage stems from Quest-GNN’s question-guided attention mechanism and multi-level message passing, which significantly reduce the impact of noise and fuse the multi-information levels comprehensively.
4.4. Efficiency
We evaluate the efficiency of retrieval across different methods, which is summarized in Table 4. Iterative retrieval based methods (Auto-RAG and KiRAG) show high efficiency because they have no extra cost of indexing or searching on graph. Graph search based methods (HippoRAG2 and RAPTOR) exhibit higher time cost because the graph search strategies are expensive. GNN based methods (Quest-GNN, G-retriever and GFM-RAG) cost less time since message passing is an efficient way to aggregate graph information  (Hamilton et al., 2017). Quest-GNN shows more retrieval time than GFM-RAG and G-retriever due to the complexity of Multi-L KG.
Table 4. Retrieval efficiency on MuSiQue. Time means the average retrieval time of each question.
Method Recall@5 Time(s)
NV-Embed-v2 68.19 0.036
Auto-RAG 61.40 0.056
KiRAG 63.05 0.072
G-retriever 66.46 0.051
GFM-RAG 67.19 0.086
RAPTOR 69.55 0.376
HippoRAG2 72.29 0.317
Quest-GNN 75.23 0.118
Table 5. Ablation study. “chunk” means performing QA task with chunk. “w/o entity”, “w/o chunk” and “w/o doc” mean no entity, chunk or document in Multi-L KG. “w/o intra” and “w/o inter” mean no intra-level message passing or no inter-level message passing. “w/o q_attn” means no question alignment. “w/o pt” means no pre-training. “w/o ft” means no fine-tuning.
@@ 4.4. Efficiency
Quest-GNN 78.54 76.02 64.65 52.64 47.47 19.03 44.63 37.85 12.50
Quest-GNN(chunk) - - - 40.18 37.55 13.26 30.25 29.41 8.71
Quest-GNN(w/o entity) 60.34 43.31 25.02 27.25 16.56 2.05 19.58 9.65 1.13
Quest-GNN(w/o chunk) 77.65 74.99 62.60 51.18 45.82 16.55 42.74 35.69 9.08
Quest-GNN(w/o doc) - - - 36.54 32.49 8.01 26.19 26.57 5.02
Quest-GNN(w/o intra) 71.33 68.42 43.15 39.88 33.39 6.93 28.76 22.85 2.61
Quest-GNN(w/o inter) - - - 40.27 35.98 4.24 28.36 19.19 2.72
Quest-GNN(w/o q_attn) 76.26 71.51 53.74 48.02 40.54 9.65 40.24 30.29 6.56
Quest-GNN(w/o pt) 63.66 53.93 36.83 35.64 28.14 5.15 24.62 19.28 2.53
Quest-GNN(w/o ft) 77.19 74.69 62.57 51.08 46.42 17.12 43.58 34.23 10.48
Table 6. The influence of GNN layer number.
MuSiQue(Rec@5) MuSiQue(F1)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24
Quest-GNN(1 layer) 73.81 69.14 55.13 46.25 41.86 11.46
Quest-GNN(2 layers) 78.54 76.02 64.65 52.64 47.47 19.03
Quest-GNN(3 layers) 76.35 74.55 61.66 51.34 45.24 16.45
Quest-GNN(4 layers) 62.60 58.29 35.37 38.93 29.15 4.61
Table 7. The influence of negative sampling number.
MuSiQue(Rec@5) MuSiQue(F1)
Method 2-hop 3-hop 4-hop 2-hop 3-hop 4-hop
HippoRAG2 79.89 73.96 48.32 53.01 44.17 10.24
Quest-GNN(Neg_N=1) 68.67 60.27 50.82 43.72 38.25 5.18
Quest-GNN(Neg_N=10) 77.97 76.16 62.14 51.03 46.68 17.62
Quest-GNN(Neg_N=30) 78.54 76.02 64.65 52.64 47.47 19.03
Quest-GNN(Neg_N=50) 77.18 76.69 63.51 51.34 46.99 18.36
Quest-GNN(Neg_N=100) 77.09 76.16 63.85 52.95 46.37 18.22
Figure 3. Influence of pre-training scale on various model dimension.
4.5. Ablation Study
We present the ablation study of Quest-GNN on Table  5. As for the chunk retrieval (Quest-GNN + chunk, w/o inter or w/o doc), we retrieve the top 10 most relevant chunks for QA and we only report the EM and F1 score since we have no ground truth for chunk retrieval. Our key findings are:
  • Quest-GNN(chunk) does not perform as well as Quest-GNN. It can be attributed to two factors: i) Quest-GNN is not directly trained on chunk labels. ii) Document generally contains more information than chunk.
  • In terms of the node in Multi-L KG, the model’s performance drops significantly without entities, as they form the basic semantic relationships in Multi-L KG. Removing entities causes substantial knowledge loss. Removing chunk nodes shows relatively less impact, as document nodes can provide global contextual information to compensate for their absence. No document lacks global context information, so it does not perform as well as Quest-GNN(chunk).
  • In terms of the message passing of Quest-GNN, no intra-level message passing severely deteriorates performance because Quest-GNN loses the ability to directly capture the structure of Multi-L KG but have to be bridged by chunk nodes. No inter-level message passing will not update the representations thus have no difference with the embeddings baselines, whose performance decrease a lot when hop number increases. The absence of question alignment has a particularly negative effect on high-hop questions, where the influence of noise will be amplified as the hop number increases.
  • In terms of the training strategy, Quest-GNN(w/o ft) can achieve competitive performance in 2-hop questions but fail to perform as well as Quest-GNN, since we pre-train Quest-GNN only on 1,2-hop questions. Quest-GNN(w/o pt) struggles to perform well because it’s likely to overfit with small number of training pairs.
4.6. Hyperparameter Discussion
We discuss some of important hyper-parameters, including the receptive field (Quest-GNN layer number), negative sampling number, pre-training data scale and information bottleneck dimension.
4.6.1. Receptive Field Discussion
We discuss the relationships between GNN layers and the multi-hop performance, the results are shown in Table  6. We find that 1-layer Quest-GNN (one intra-level + one inter-level) have limited performance and the gap between 2-layer Quest-GNN becomes bigger as the hop number increase, it is because 1-layer Quest-GNN can only aggregate 2-hop information. The receptive field limits the potential of Quest-GNN. The 2-layer Quest-GNN achieves the best performance among 2,3,4 hop questions because it can aggregate more useful information. However, as the layer number increases (4 layers), the performance will suffer a lot because of the over-smoothing problem of GNN  (Rusch et al., 2023; Chen et al., 2020a).
4.6.2. Negative Sampling Number
Table 7 shows the impact of negative sampling number on model’s performance. According to the results, we find that low sampling number (Neg_N = 1) has limited performance due to inadequate discrimination between target documents and distractions. High sampling number (Neg_N=50, 100) shows limited benefits, because excessive easy negatives fail to provide meaningful guidance for Quest-GNN training.
4.6.3. Pre-training Scale and Information Dimension
We pre-train Quest-GNN on the sythesized QA pairs from 0 to 150k across information dimension from 32 to 512 respectively, the results are shown in Figure  3. We conduct experiments on the MuSiQue dataset. As for the pre-training scale, the results show that insufficient pre-training data leads to sub-optimal performance across all the dimensions. As the amount of pre-training data increases, the performance of Quest-GNN gets better, although the marginal improvement decreases. As for the information dimension, the results indicate that increasing the information dimension improves the performance of Quest-GNN, as larger dimensions enhance model expressiveness through increased model scale. This observation is consistent with the model scaling analysis presented in GFM-RAG  (Luo et al., 2025). We believe that further improvements can be achieved by combining a larger model with a substantial amount of pre-training data. In this paper, we set the dimension as 128 considering the balance between performance and computational limitation.
5. Conclusion
@@
In this paper, we propose a novel graph learning method for multi-hop questions retrieval. We first design a Multi-information Level Knowledge Graph(Multi-L KG) to model the multi-granular relationships among documents. Then we introduce Question-Adaptive Graph Neural Network(Quest-GNN). The Quest-GNN aggregates information from multi-levels, together with question alignment, to ensure comprehensive question specific representations while filtering out the noise. The pre-training strategy further enhance the ability of Quest-GNN. Extensive experiments shows the effectiveness of our method in multi-hop questions.
@@ Appendix A Appendix
A.1. Dataset Settings
Table 8. Dataset statistics
@@ A.1. Dataset Settings
The statistics of benchmark datasets are shown in Table  8. For the MuSiQue, 2Wiki and HotpotQA, the extra sample have at least 60%, 60% and 50% supported documents in existing corpora.
A.2. Baseline Settings
BM25 is implemented as BM25S (Lù, 2024). As for the text embedding based models, we use the NV-Embed-v2-7B and GTE-Qwen2-7B-Instruct from Huggingface  (Wolf et al., 2019). For the all the other baselines we use their official implementations. For iterative retrievers, we set the max retrieval number as 5. For GFM-RAG, we use the official settings to pre-train and fine-tune on our corpora. For GNN-RAG we use ReaRev  (Mavromatis and Karypis, 2022) as GNN reasoning. All other baseline settings are as their official suggestions.
A.3. QA Performance on GPT-4o-mini
@@
Table 9. QA performance on GPT-4o-mini. We highlight the best with bold and the second best with under line.
MuSiQue 2Wiki HotpotQA
Method EM F1 EM F1 EM F1
No Retrieval 10.81 21.39 31.83 36.51 27.44 41.47
BM25 12.54 23.88 32.25 40.30 34.50 52.62
Contriever+IRCoT 13.90 25.28 29.44 34.66 30.17 45.55
GTE-Qwen2-7B 31.15 39.88 50.75 52.37 52.19 66.17
NV-Embed-v2 28.07 37.15 49.22 51.51 54.85 68.43
GNN-RAG 23.02 30.12 34.52 48.25 49.12 58.99
GFM-RAG 28.30 36.81 43.79 53.26 51.86 64.73
RAPTOR 32.93 41.04 45.43 56.24 55.19 68.16
GraphRAG 33.31 40.13 50.99 58.08 56.45 69.24
HippoRAG2 33.12 40.30 51.49 60.18 56.39 70.53
Quest-GNN(Ours) 34.27 41.78 53.52 61.63 56.04 69.73
@@
Table 9 presents QA performance using GPT-4o-mini. The results show that Quest-GNN achieves competitive EM and F1 scores, demonstrating its compatibility with different LLM backbones. Notably, all methods show reduced performance compared to Llama-3-70B-Instruct, primarily due to the intrinsic knowledge gap between GPT-4o-mini and Llama-3.3-70B-Instruct. Other observations align with the report of Table  2.
A.4. Case Study and Discussion